Cardanos method for finding roots. x + 6x²+68-8=0
Answer #1


30, S=31K+124R7 -3f2+252 - T= 31 Rr SQR4R2 2 282-252 SATE 22-4 2/2 + 3/2-22 S-T = 32+252 -352-252 x3 +68² +67-8=0 comparing w